OpenStudy (anonymous):

HELP ME PLEASE ! 1. What is the solution of n² – 49 = 0? (1 point) –7 7 ±7 no solution 2. What is the solution of x² + 64 = 0? (1 point) –5 8 ±8 no solution 3. What is the side length of a square with an area of 144x²? (1 point) 12 12x ±12x no solution 5. What is the value of z so that –9 and 9 are both solutions of x² + z = 103? (1 point) –22 3 22 184

OpenStudy (kainui):

These appear to be copy and pasted, why don't you describe to me what your problem is in solving these?

OpenStudy (anonymous):

I know the first answer but I just don't understand how to get the answers. can u please explain the steps someone said u add the square root to the 49 but i get like 2. something if I do that

OpenStudy (kainui):

Ok, so if you look at: n² – 49 = 0 we can rearrange it by adding 49 to both sides, since that will still keep it equal on both sides, right? n²=49 Now we can say, hey, what multiplied by itself is 49? I suggest you rearrange most of the problems similarly to this, and keep in mind that a negative number times a negative number is a positive number.

OpenStudy (anonymous):

So it would look like this 49^2=49?

OpenStudy (anonymous):

or 49^2-49=0?

OpenStudy (anonymous):

I got 2.35?

OpenStudy (kainui):

No, it will look like this: n² – 49 = 0 add 49 to both sides, since changing both sides by the same amount keeps it equal. n² – 49 +49= 0+49 Now you can simplify it to n² =49 and the take the square root of both sides n=7 and n=-7 Show me what you're doing to get 2.35, because that is just plain wrong and I want to help you figure out where you need to fix it so you can get it right. =)

OpenStudy (anonymous):

Oh okay I understand it now thank you! does square root mean basically 7*7 cuz that equals 49? I was doing it in the calculator wrong that is why I got 2.35

OpenStudy (anonymous):

so the next one would look like this x^2+64=0 then x^2+64+64=0+64 then simplify x^2=64 =8 =)

OpenStudy (anonymous):

You square root it to get rid of the n^2. and when solving equations what you do to one side you must apply to both sides. the sqrt of 49 is 7 because 7x7=49. :)

OpenStudy (anonymous):

and yep, you got that last one correct! good job.

OpenStudy (stormrider4):

1 is b 2 is d 3 is a